OpenStudy (anonymous):

the perimeter of a rectangle is 38.6m the width of the rectangle is 7.2 what is the lengh

OpenStudy (texaschic101):

formula for perimeter of a rectangle : P = 2L + 2W P = 38.6 W = 7.2 now we sub 38.6 = 2L + 2(7.2) 38.6 = 2L + 14.4 38.6 - 14.4 = 2L 24.2 = 2L 24.2/2 = L 12.1 = L check.. P = 2L + 2W 38.6 = 2(12.1) + 2(7.2) 38.6 = 24.2 + 14.4 38.6 = 38.6 (correct) Length (L) = 12.1